OpenStudy (anonymous):

3 – 7 • 5 + 2 • 4 A. –40 B. –24 C. –12 D. 28

OpenStudy (anonymous):

I think the answer is a

OpenStudy (secret-ninja):

Use PEMDAS to solve P- PARENTHESIS E- EXPONENTS M- MULTIPLICATION D- DIVISION A- ADDITION S- SUBTRACTION

OpenStudy (secret-ninja):

3 - 7 x 5 + 2 x 4 7x5=35 2x4=8 3-35+8 35+8=43 3-43 = -40 A is correct.

OpenStudy (secret-ninja):

Good job. ^-^

OpenStudy (anonymous):

I did that

OpenStudy (secret-ninja):

Click the "Close" button when you're done. :)

OpenStudy (anonymous):

i know. i added you to a new post